OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!! A cone has a diameter of 24 in. and a slant height of 15 in.What is the approximate volume of the cone? Use 3.14 to approximate the value of pi. Round your answer to the nearest tenth if necessary.

OpenStudy (anonymous):

Formula is \(\dfrac{1}{3}\) x \(\Large\pi\) x radius^2 x height

OpenStudy (anonymous):

So can you figure out \(\dfrac{1}{3}\) x 3.14 x 12^2 x height? (Radius us 12 because it is half the diameter)

OpenStudy (anonymous):

I know...What I am confused about is how to work it... I will post the picture

OpenStudy (anonymous):

We don't need the picture.

OpenStudy (anonymous):

Ok, since the radius is half the diameter, what is 12^2?

OpenStudy (anonymous):

144

OpenStudy (anonymous):

how would you figure out the height?

OpenStudy (anonymous):

@answermyquestions

OpenStudy (anonymous):

It says the height is 15.